OpenStudy (anonymous):

I really need help figuring out probability, for example... USA Today reported that approximately 25% of all state prison inmates released on parole become repeat offenders while on parole. Suppose the parole board is examining five prisoners up for parole. Let x = number of prisoners out of five on parole who become repeat offenders. 0=.209 1=.377 2=.227 3=1.49 4=.037 5=.001 Find the probability that one or more of the five parolees will be repeat offenders. (Round your answer to three decimal places.)

OpenStudy (perl):

Let X = # of parolees out of 5 prisoners paroled who become repeat offenders. you want P(X>= 1)

OpenStudy (perl):

it is easier to do the complement. P( A) = 1 - P ( A ' )

OpenStudy (anonymous):

...still don't know how to figure this out

OpenStudy (anonymous):

what goes where

OpenStudy (perl):

P( X >= 1 ) = 1 - P ( X < 1 ) = 1 - P(x= 0)

OpenStudy (perl):

do you agree so far?

OpenStudy (anonymous):

yes

OpenStudy (perl):

i am just planning out the theory, then i will plug in formula

OpenStudy (perl):

ok we want the probability of F F F F F (or S S S S S depending how you look at it )

OpenStudy (perl):

F = fail, S = success

OpenStudy (perl):

so we want ( prob. did not repeat offend) x ( prob. did not repeat offend)x( prob. did not repeat offend) x ( prob. did not repeat offend) x ( prob. did not repeat offend)

OpenStudy (anonymous):

still lost